# OWLed: Outlier-weighed Layerwise Pruning for Efficient Autonomous Driving Framework

### v2

@@ IV-C Ablation study: impact of pruning model encoder
Our model consists of two main components: an encoder and LLaMA-7B as the LLM. The encoder has 9053440 parameters, while LLaMA-7B has 6738415616 parameters. The encoder’s parameter size is approximately three orders of magnitude smaller than the LLM’s. Given this significant size difference, we conduct an ablation study to investigate how pruning the relatively compact encoder affects overall model performance.
@@
Sparsity_ratio Method GPT Grading $\uparrow$
@@
TABLE III: GPT grading of pruning encoder ablation.
@@
TABLE IV: Evaluation for calibration samples ablation.
@@
As shown in Table II, at sparsity levels of 0.3 and 0.4, OWLed outperforms both OWLed-separate and OWLed-global across multiple evaluation metrics. The results in Table III are consistent with the trends observed in Table II. This indicates that pruning the encoder has a significant impact on model performance when pruning the same number of parameters. A possible explanation is that the encoder, despite its smaller size, plays a crucial role in processing input data, which further influences the overall performance. Therefore, in this scenario, we do not recommend pruning the encoder. Notably, since the encoder’s parameter number is much less than that of the LLM, focusing on pruning the LLM may lead to better model performance.
@@ IV-D Ablation study: number of calibration samples
[…]
As shown in Table IV, using 32 calibration samples fails to achieve optimal performance across all tasks. For sample sizes exceeding 32, we observe that different sets only attain superior performance on specific tasks or metrics. This observation indicates that while increasing the sample size beyond 32 can enhance performance in certain areas, it doesn’t yield consistent, comprehensive improvements. Therefore, we recommend selecting a moderate number of calibration samples to reduce computational costs and accelerate the pruning process.
Table V shows that while 32 calibration samples yield the lowest GPT grading score, the scores for larger sample sizes are comparatively similar. This aligns with our observations from Table IV: Beyond the 32-sample threshold, increasing the number of calibration samples does not lead to substantial improvements in model performance. The consistency between these two evaluation methods strengthens our recommendation to use a moderate number of calibration samples. This approach effectively balances model performance with computational efficiency, optimizing the calibration process without incurring unnecessary computational overhead.
@@
TABLE V: GPT Grading for calibration samples ablation.
@@ V Conclusion
This paper presents OWLed, an efficient autonomous driving framework that employs outlier-weighted layerwise pruning to compress LLM without fine-tuning. By assigning non-uniform sparsity ratios based on the distribution of outlier features and incorporating driving environment data into calibration and pruning, our method significantly reduces computational demands while maintaining performance. We also found that the encoder is more sensitive to pruning than the LLM. Our experiments show that OWLed outperforms existing methods in perception, action prediction, and language understanding. These findings demonstrate the potential of combining advanced pruning techniques with LLMs to develop efficient and robust autonomous driving systems capable of handling complex scenarios. Future work includes applying weight-low-rank and quantization techniques to further increase model efficiency.

### v3

@@ IV-C Ablation study: impact of pruning model encoder
Our model consists of two main components: an encoder and LLaMA-7B as the LLM. The encoder has 9.1e+6 parameters, while LLaMA-7B has 6.7e+9 parameters. The encoder’s parameter size is approximately three orders of magnitude smaller than the LLM’s. Given this significant size difference, we conduct an ablation study to investigate how pruning the relatively compact encoder affects overall model performance.
@@
Sparsity ratio Method GPT Grading $\uparrow$
@@
TABLE III: GPT grading results of language understanding capabilities across OWLed variants at 0.3 and 0.4 sparsity ratios, respectively.
@@
TABLE IV: Evaluation for calibration samples ablation. Performance comparison across different numbers of calibration samples (32 to 512) on various metrics.
@@
As shown in Table II, at sparsity levels of 0.3 and 0.4, OWLed outperforms both OWLed-separate and OWLed-global across multiple evaluation metrics. The results in Table III are consistent with the trends observed in Table II. This indicates that pruning the encoder has a significant impact on model performance when pruning the same number of parameters.
A possible explanation for the encoder’s sensitivity to pruning is because of its role in translating sensor data into compact vector representations. Even minor distortions in these representations can propagate through the LLM, significantly impacting decision-making. In contrast, the LLM primarily processes high-level features and is acceptable to reduce part of the parameter due to its massive parameter count (6.7B vs. encoder’s 9.1M parameters).
Therefore, in this scenario, we do not recommend pruning the encoder. Notably, since the encoder’s parameter number is much less than that of the LLM, focusing on pruning the LLM may lead to better model performance.
@@ IV-D Ablation study: number of calibration samples
The calibration strategy consists of calibration data type and calibration data number. An ablation study is conducted to investigate the impact of the number of calibration data on the OWLed performance. We vary the number of calibration data from the set (32, 64, 128, 256, 512)($\text${32, 64, 128, 256, 512}) and present the results in Tables IV and V. We set the sparsity ratio as 0.4.
As shown in Table IV, using only 32 calibration samples results in suboptimal performance across all metrics, this could because insufficient samples fail to capture the full distribution of feature importance across different driving scenarios. Although increasing samples beyond 32 improves performance, the gains become marginal after 128 samples. We recommend using 256 calibration samples in this task as an optimal balance between performance and computational efficiency.
Table V shows that while 32 calibration samples yield the lowest GPT grading score, the scores for larger sample sizes are comparatively similar. This aligns with our observations from Table IV: Beyond the threshold of 128 samples, increasing the number of calibration samples does not lead to a significant improvement in model performance. The consistency between the two evaluation methods strengthens our recommendation to select a moderate number of calibration samples, rather than simply increasing the number of samples. This approach effectively balances model performance with computational efficiency, optimizing the calibration process without incurring unnecessary computational overhead.
@@
TABLE V: GPT grading results of calibration samples ablation. Analysis of model response quality across different calibration sample sizes (32 to 512).
@@ V Conclusion
This paper presents OWLed, an efficient autonomous driving framework that employs outlier-weighed layerwise pruning to compress LLM without fine-tuning. By assigning non-uniform sparsity ratios based on the distribution of outlier features and incorporating driving environment data into calibration and pruning, our method significantly reduces computational demands while maintaining performance. We also found that the encoder is more sensitive to pruning than the LLM. Our experiments show that OWLed outperforms existing methods in perception, action prediction, and language understanding. These findings demonstrate the potential of combining advanced pruning techniques with LLMs to develop efficient and robust autonomous driving systems capable of handling complex scenarios. The OWLed’s effectiveness may be constrained in scenarios with highly dynamic environments or unseen driving conditions, where the pre-computed outlier distributions might not fully capture the complexity of the situation. Future work includes applying weight-low-rank and quantization techniques to further increase model efficiency.